1. Find the equation of the curve at every point that passes through the point (0,-1) and which the normal line at any point (x,y) has a slope of – 1/2(x+1). Ans. y=x2+2x-1
  2. Given the family of curves x2+y2=cx, find the family of curves of the orthogonal trajectories. Ans. x2+y2=ky
1
Expert's answer
2021-11-19T15:19:34-0500

1.

"slope_1=m_1=-\dfrac{1}{2x+1}"

Then the tangent line at any point"(x,y)" has a slope  


"slope_2=m_2=-\dfrac{1}{m_1}=2x+1"

"y'=2x+1"

Integrate


"y=\int(2x+1)dx=x^2+x+C"

The curve passes through the point "(0,-1)"


"-1=(0)^2+0+C=>C=-1"

The equation of the curve is

"y=x^2+x-1"

2.  Given family of curves is "x^2+y^2=cx"

Differentiate both sides with respect to "x"


"2x+2yy'=c"

"y'=\dfrac{c-2x}{2y}"

"y'=\dfrac{x^2+y^2-2x^2}{2xy}"

"y'=\dfrac{y^2-x^2}{2xy}"

If any curve intersects orthogonally at point "(x,y)," then (if its slope is "y'") we must have


"\dfrac{dx}{dy}=-\dfrac{y^2-x^2}{2xy}=\dfrac{x}{2y}-\dfrac{y}{2x}"

Solving the above differential equation, we get


"x=ty"

"\dfrac{dx}{dy}=t+y\dfrac{dt}{dy}"

"t+y\dfrac{dt}{dy}=\dfrac{t}{2}-\dfrac{1}{2t}"

"y\dfrac{dt}{dy}=-\dfrac{t}{2}-\dfrac{1}{2t}"

"t\dfrac{dt}{t^2+1}=-\dfrac{dy}{2y}"

Integrate


"\int t\dfrac{dt}{t^2+1}=-\int\dfrac{dy}{2y}"

"t^2+1=\dfrac{k}{y}"

"(\dfrac{x}{y})^2+1=\dfrac{k}{y}"

"x^2+y^2=ky"